# From Lived Experience to Insight: Unpacking the Psychological Risks of Using AI Conversational Agents

### v1

@@ 3.1.1. Structuring psychological risks
  • Context: The context in which a user engages with an AI conversational agent encompasses the surrounding conditions, user-specific circumstances, and additional relevant information that collectively influence the interaction between the individual and the AI conversational agent. This context includes, but is not limited to, the individual’s background (such as cultural, religious, and demographic attributes) mental and physical health status, intent, as well as external factors (such as external environment).
@@
  • AI Behavior: The behavior of an AI conversational agent (such as ChatGPT, Microsoft Copilot etc.) refers to the actions performed by the AI conversational agents. This behavior encompasses several aspects of agent responses such as (but not limited to) content, tone, voice, choice of words, language, punctuation, obedience to user inputs, adaptability, information sharing, etc. Furthermore, these behaviors can be shown in various modalities such as (but not limited to) text, images, voice, videos, etc.
@@
  • Negative Psychological Impact: The AI conversational agent’s action creates a risk of negative psychological and/or social impact to one or more individuals. Negative psychological impact in this case refers to any negative impact on an individual’s mental or emotional well-being, which can manifest as exacerbation of mental health condition, reduced self-esteem, or other issues such as physiological harm. These negative impacts also encompass the impacts on an individual’s social interactions, relationships, and standing within a community, potentially leading to social isolation, stigmatization, or discrimination.
@@
Throughout our study, we leverage this structure to collect data around negative psychological experiences with AI. Of these five components, we largely focus on context, AI behavior, and negative psychological impact because likelihood and temporality can be highly idiosyncratic.
Category Details Count of Participants (w.r.t N=279)
Submission Source UserTesting
Prolific
Social Media
268 ($\sim$96.1%)
8 ($\sim$2.9%)
3 ($\sim$1.1%)
Age 18–25 years
26–35 years
36–45 years
46–55 years
56–65 years
Preferred not to disclose
114 ($\sim$40.9%)
100 ($\sim$35.8%)
52 ($\sim$18.6%)
10 ($\sim$3.6%)
2 ($\sim$0.7%)
1 ($\sim$0.4%)
Gender Woman
Man
Non-binary/gender diverse
Preferred not to disclose
145 ($\sim$52.0%)
123 ($\sim$44.1%)
10 ($\sim$3.6%)
1 ($\sim$0.4%)
Primary Language English
Other (Spanish, Mandarin Chinese, Vietnamese, Farsi, Portuguese)
271 ($\sim$97.1%)
8 ($\sim$2.9%)
Familiarity with AI Agents Interacted for 1+ years
Interacted for 6+ months
Interacted for ¡6 months
165 ($\sim$59.1%)
88 ($\sim$31.5%)
26 ($\sim$9.3%)
Frequency of AI Agent Use Once or more per day
Once or more per week
Once or more per month
Less than once per month
144 ($\sim$51.6%)
104 ($\sim$37.3%)
26 ($\sim$9.3%)
5 ($\sim$1.8%)
Table 1. Participants’ Data Overview
3.1.2. Recruitment
Our primary method of recruitment was through UserTesting, an online platform used for gathering public feedback on product and services. On the platform, we screened participants for several inclusion criteria: (1) having prior experience using AI conversational agents, (2) having experienced a negative psychological impact from using AI conversational agents, (3) self-identifying as a person with lived experience in mental health, and (4) 18 years or older residing in the US who comprehended English. To expand our recruitment, we also used social media platforms and Prolific. We posted the link to our study on the study team’s personal Twitter and LinkedIn accounts. We were permitted to post the study on 4 subreddit groups (r/ChatGPT, r/Bard, r/SampleSize, r/SurveyExchange). We prematurely terminated our study task on Prolific because most submissions failed to meet the inclusion criteria despite using platform filters and providing explicit instructions. The survey study ran from August to September of 2024. UserTesting and Prolific participants were compensated about US \$10, and US \$4 respectively, after platform service fees. We did not compensate participants recruited from social media.
3.1.3. Participant demographics and data
Given that the survey format allowed each participant to provide up to three past lived experiences, we received a total of 297 scenarios from 283 participants. We excluded 7 scenarios from 4 participants because there were incomprehensible or were not negative psychological experiences, leaving us with 290 scenarios from 279 participants. We received 268 submissions from UserTesting, 8 from Prolific, and 3 from social media.
114 (40.9%) participants were aged 18-25 years, 100 (35.8%) were 26–35 years old, 52 (18.6%) were 36–45 years, 10 (3.6%) were 46–55 years, 2 (0.7%) were 56-65 years, and 1 preferred not to disclose age. 145 (52.0%) participants identified as woman, 123 (44.1%) as man, 10 (3.6%) as non-binary/gender diverse, and 1 preferred not to disclose gender identity. 271 (97.1%) participants responded that they typically think in English. Other languages included Spanish, Mandarin Chinese, Vietnamese, Farsi, and Portuguese.
On a 5-point scale (1: not at all, 5: very familiar), the average familiarity with conversational AI agents was 3.93 ($\sigma$=0.71). 165 (59.1%) participants interacted with AI agents for 1 or more years, 88 (31.5%) for 6 or more months, and 26 (9.3%) for less than 6 months. 144 (51.6%) used AI agents once or more per day, 104 (37.3%) used once or more per week, 26 (9.3%) used once or more per month, and 5 used less than once per month. The top 3 most frequent purposes for using AI conversational agents were Researching (86.4%), Getting advice (76.3%), and Learning (73.8%). On a 5-point scale (1: not at all, 3: somewhat interested, 5: very interested), the average interest level in using AI conversational agents for mental health support was was 3.26 ($\sigma$=1.30). Specifically, 46.2% of the participants reported to being moderately interested (4) or very interested (5). Table 1 outlines the characteristics of our survey participants.
3.1.4. Survey design
After acknowledging consent to participate in the study, participants were asked a series of questions aimed at understanding the participant, their AI experience, and their negative psychological experience. The survey was divided into the following sections:
@@ 3.1.4. Survey design
  • Context: In the context specific section, we asked participants to share any specific aspects about themselves or situation that might help us understand the experience. We then asked how long ago the experience took place, which AI conversational agent they were using, interaction modalities (e.g., text, voice, image, video) and language, and the purpose of engagement.

### v3

@@ A.1. Phase 1: Structuring psychological risks
  • Context: The context in which a user engages with an AI conversational agent encompasses the surrounding conditions, user-specific circumstances, and additional relevant information that collectively influence the interaction between the individual and the AI conversational agent. This context includes, but is not limited to, the individual’s background (such as cultural, religious, and demographic attributes), mental and physical health status, intent, as well as external factors (such as the external environment).
@@
  • AI Behavior: The behavior of an AI conversational agent (such as ChatGPT, Microsoft Copilot, etc.) refers to the actions performed by the AI conversational agents. This behavior encompasses several aspects of agent responses such as (but not limited to) content, tone, voice, choice of words, language, punctuation, obedience to user inputs, adaptability, information sharing, etc. Furthermore, these behaviors can be shown in various modalities such as (but not limited to) text, images, voice, videos, etc.
@@
  • Negative Psychological Impact: The AI conversational agent’s action creates a risk of negative psychological and/or social impact to one or more individuals. Negative psychological impact in this case refers to any negative impact on an individual’s mental or emotional well-being, which can manifest as exacerbation of mental health conditions, reduced self-esteem, or other issues such as physiological harm. These negative impacts also encompass the impacts on an individual’s social interactions, relationships, and standing within a community, potentially leading to social isolation, stigmatization, or discrimination.
@@
A.2. Phase 1: Survey design
Consented participants were asked a series of questions aimed at understanding the participant, their AI experience, and their negative psychological experience: The survey was divided into the following sections:
@@ A.2. Phase 1: Survey design
  • Context: In the context-specific section, we asked participants to share any specific aspects about themselves or their situation that might help us understand the experience. We then asked how long ago the experience took place, which AI conversational agent they were using, interaction modalities (e.g., text, voice, image, video) and language, and the purpose of engagement.